OpenStudy (anonymous):

Last problem I need help with. A tugboat goes 180 miles upsteam in 15 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current.

OpenStudy (lgbasallote):

upstream speed = speed of boat - speed of current downstream speed = speed of boat + speed of current upstream speed is \[\frac{180}{15} \implies 12 mph\] downstream speed is \[\frac{180}{5} \implies 36 mph\] so we use the formula i gave upstream = b - c downstream = b +c where b = speed of boat c = speed of current 12 = b - c 36 = b +c ======= 48 = 2b <---add the two equations now solve for b 2b = 48 b = 48/2 b = 24 mph since downstream speed is 36... dwnstream = b +c 36 = 12 + c 36 - 12 = c 24 = c

OpenStudy (lgbasallote):

uhh i made a mistake.. downstream = b + c 36 = 24 + c 36 - 24 = c 12 = c

OpenStudy (lgbasallote):

did you get that @ldepallo

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so which one is the speed of the tugboat and which one is the current?

OpenStudy (lgbasallote):

look at my solution. i mentioned it *if* you read the solution

OpenStudy (anonymous):

ok